OpenStudy (anonymous):

rewrite (5 sqrt x) /x

OpenStudy (anonymous):

\[(5\sqrt{x} ) / x\]

OpenStudy (anonymous):

you can rewrite it as \(5\sqrt{x}\) if you like, because \(\frac{\sqrt{x}}{x}=\sqrt{x}\)

OpenStudy (anonymous):